kekeman:

For what value of θ is cscθ undefined? A.) 0 degrees B.) 45 degrees C.) 90 degrees D.) 270 degrees

surjithayer:

\[\sin \theta=0,\theta=0,180,360\] \[\csc \theta ~is~undefined~at~these~points.\] so A